If the hypotenuse of a right triangle is 8 feet less than 3 times shorter leg and the longer leg is 8 feet more than 2 times the shorter leg, what is the length of all three sides?
The answer is one of these:
a) 15ft, 36ft, 39ft
b) 18ft, 44ft, 46ft
c) 20ft, 48ft, 52ft
d) 30ft, 68ft, 82ft

Answer by scott8148(6628)

by Pythagoras ___ x^2 + (2x+8)^2 = (3x-8)^2

x^2 + 4x^2 + 32x + 64 = 9x^2 - 48x + 64

0 = 4x^2 - 80x

0 = x^2 - 20x

0 = x(x - 20)

x = 0

x - 20 = 0 ___ x = 20